# MOKA: Open-Vocabulary Robotic Manipulation through Mark-Based Visual Prompting

### v1

@@ Abstract
Open-vocabulary generalization requires robotic systems to perform tasks involving complex and diverse environments and task goals. While the recent advances in vision language models (VLMs) present unprecedented opportunities to solve unseen problems, how to utilize their emergent capabilities to control robots in the physical world remains an open question. In this paper, we present Marking Open-vocabulary Keypoint Affordances (MOKA), an approach that employs VLMs to solve robotic manipulation tasks specified by free-form language descriptions. At the heart of our approach is a compact point-based representation of affordance and motion that bridges the VLM’s predictions on RGB images and the robot’s motions in the physical world. By prompting a VLM pre-trained on Internet-scale data, our approach predicts the affordances and generates the corresponding motions by leveraging the concept understanding and commonsense knowledge from broad sources. To scaffold the VLM’s reasoning in zero-shot, we propose a visual prompting technique that annotates marks on the images, converting the prediction of keypoints and waypoints into a series of visual question answering problems that are feasible for the VLM to solve. Using the robot experiences collected in this way, we further investigate ways to bootstrap the performance through in-context learning and policy distillation. We evaluate and analyze MOKA’s performance on a variety of manipulation tasks specified by free-form language descriptions, such as tool use, deformable body manipulation, and object rearrangement.
Figure 1: To solve manipulation tasks with unseen objects and goals, MOKA employs a VLM to generate motions through a point-based affordance representation (plotted as colorful dots on the images). By annotating marks (e.g., candidate points, grids, and captions) on the observed 2D image, MOKA converts the motion generation problem into a series of visual question-answering problems that the VLM can solve.
@@ I Introduction
The pursuit of open-vocabulary generalization poses a major challenge for robotic systems: Solving tasks in unseen environments given new user commands necessitate methods that can deal with the vast diversity and complexity of the physical world. An appealing prospect for handling this challenge is to employ large pretrained models by encapsulating extensive prior knowledge from broad data and bringing it to bear on novel problems. Recent advances in large language models (LLMs) and vision-language models (VLMs) provide particularly promising tools in this regard, with their emergent and fast-growing conceptual understanding, commonsense knowledge, and reasoning abilities [8, 39, 40, 4, 7, 1, 41, 22, 42, 23]. However, existing large models pre-trained on Internet-scale data still lack the capabilities to understand 3D space, contact physics, and robotic control, not to mention the knowledge about the embodiment and environment dynamics in each specific scenario, creating a large gap between the promising trend in computer vision and natural language processing and applying them to robotics. It remains an open question how such tools can guide a robotic system to solve manipulation tasks by interacting with the physical world.
@@
Recently, a growing body of research has been dedicated to utilizing pre-trained large-scale models for robotic control. By incorporating broad knowledge, these approaches directly prompt or fine-tune the large models to generate plans [2, 16, 15, 5], rewards [28, 20, 30, 56], codes [25, 45, 49], etc. Despite the encouraging results they have demonstrated, these approaches are subject to notable limitations. Since the advances in LLMs precede VLMs, many previous approaches first process the raw sensory inputs to obtain the language description of the environment and then query LLMs to perform reasoning and planning in the language domain. However, relying solely on high-level language descriptions may overlook the nuanced visual details of environments and objects, which are vital for accurately completing tasks. In addition, existing approaches usually require non-trivial effort in designing in-context examples [17, 25, 45] to ensure LLMs can produce desired predictions on similar tasks. As a result, the tasks that can be solved by these approaches are largely constrained by such manual efforts.
@@
In this work, we study how to effectively endow robots the ability to solve novel manipulation tasks specified by free-form language instructions using VLMs. Our key insight is to find an intermediate affordance representation that connects the VLM’s prediction on images with the robot’s motion in the physical world. This affordance representation should satisfy two critical requirements. First, it should be feasible for the VLM to predict given the visual observation of the environment and task description. Second, it should compactly capture the information that well characterizes the important properties of the robot’s motion, such that it can be easily executed on the robot.
@@
To this end, we propose Marking Open-vocabulary Keypoint Affordances (MOKA), an approach that employs VLMs for robotic manipulation through mark-based visual prompting. As shown in Fig. 1, MOKA leverages a compact affordance representation consisting of a set of keypoints and waypoints, defined on open sets of objects and tasks. This point-based affordance representation is then used to specify the desired motion for the robot to solve the task. To generate the motions given the free-form language descriptions, MOKA uses hierarchical visual prompting to convert the affordance reasoning problem into a series of visual question answering problems.
Drawing inspirations from recent advances in visual question-answering [51], we use mark-based visual prompting to enable the VLM to attend to the important visual cues in the observation image and further simplify the point generating problem into multiple choice questions. As shown in the top-left part of Fig. 1, we plot the keypoints on the image, and query the VLM to select the keypoints that result in the desired motion. The predicted keypoints and waypoints are used for specifying a trajectory through a waypoint-following motion, which can represent a wide range of manipulation skills such as picking, placing, pressing, tool-use, etc.
@@
We demonstrate the effectiveness of MOKA on a variety of manipulation tasks, with robustness across the variations of instructions, objects, and scenes. We further use MOKA to collect successful trajectories in each task. The collected trajectories can be used as in-context examples to further bootstrap the performance of VLM, and can also be leveraged as demonstrations to train a better student policy. Our experiments show that MOKA can achieve state-of-the-art performance on our proposed evaluation tasks in a zero-shot manner, with consistent improvements using clean and intuitive in-context examples. To summarize, our contributions are:
@@
  • We propose a mark-based visual prompting approach that converts the affordance reasoning problem into a series of visual question answering problems.
  • We demonstrate that the resultant approach, MOKA, can effectively generate motions to solve diverse open-vocabulary manipulation tasks.
@@ II Related Work
Figure 2: Overview of MOKA. We propose a hierarchical approach to prompt the VLM to perform affordance reasoning. On the high-level, we query the VLM to decompose the free-form language description of the task into a sequence of subtasks and summarize the subtask information. On the low-level, the VLM is prompted to produce the keypoints and additional attributes for the affordance representation defined in Sec. IV-A.
@@
Large language models and vision-language models. Recent developments in various domains and applications have been greatly influenced by the substantial progress achieved through large language models (LLMs) and vision language models (VLMs) [8, 39, 40, 4, 7, 1, 41, 22, 42, 23]. While these models can already solve various tasks in a zero-shot manner [1], well-designed prompts still serve an important role in further eliciting more advanced capabilities. As demonstrated by Brown et al. [4], few-shot prompting can match or surpass the performance of fine-tuning on LLMs. Additionally, other prompting techniques [50, 57, 19, 55] have been proposed to improve LLMs’ reasoning capabilities. Although these methodologies may initially appear enigmatic, they have been empirically validated to consistently demonstrate scalability across various models [4, 7, 1, 54]. Apart from the language prompts, recent vision models [18, 58, 21] are capable of supporting visual prompts, including points, boxes, masks, and texts. Such visual prompts exhibit greater diversity in both form and content, harnessing vision-language models for various application scenarios like perception [53], image editing [44] and reasoning [51].
@@
To build an autonomous agent capable of making decisions in an unstructured environment, the incorporation of robust visual reasoning capabilities becomes imperative. Although the current generation of VLMs cannot seamlessly engage in zero-shot reasoning, they can still be effectively harnessed across a diverse spectrum of tasks. Prior work SoM [51] draws visual marks on the objects in an image using numbers and segmentation masks, and demonstrates the mark-based visual prompting scheme unleashes the reasoning capabilities of GPT-4V, such as object counting and inferring spatial relationship. Different from SoM [51], we want to leverage GPT-4V for open-vocabulary robot manipulation. We represent each manipulation phase with a set of affordance points and motion waypoints. Instead of using object segmentation masks, we use keypoints and grid cells as visual prompts as shown in Fig. 1, and then query GPT-4V to choose the affordance points and motion waypoints from the visual marks, followed by executable point-based motion plans.
@@
Foundation models in robotics. Building upon the successes of large language models, the field of robotics is currently witnessing a rising interest in utilizing LLMs in various application scenarios. LLMs is capable of generating high-level task plan in natural language [2, 16, 15, 5], executable programs [25, 45, 49] or value function [17, 26] for low-level behaviours, environmental reward and feedback for reinforcement learning [28, 20, 30, 56]. However, these approaches necessitate the conversion of both the task and the observations into the textual form. While this can be easily accomplished in simulated environments with ground-truth object state, tasks in real world require the utilization of robust and precise perception modules. To perform open-ended navigation and manipulation in real world [46, 5, 10], open-vocabulary vision foundation models [18, 58, 21, 27, 34, 24] are often used to extract visual scene information before converting the observation to textual form. However, the process of converting an image into text may result in the loss of important details, such as shape and geometric information. With the recent advancements in vision language models (VLMs), it is now conceivable that the role of state estimation combined with language models can be replaced by a single, powerful VLM, such as GPT-4V, which is leveraged by MOKA.  Hu et al. [14] utilizes GPT-4V to get semantic language plans and convert each language plan to a set of pre-defined low-level skills. Different from Hu et al. [14], we represent manipulation affordance and motion using a set of points, and query GPT-4V to select the corresponding points (e.g. grasping point, function point) in each task, which can be directly translated to low-level point-based motions. While preserving the visual reasoning capabilities of VLMs, MOKA provides a framework with more general and flexible low-level motion, which doesn’t require prior knowledge about a specific task compared to pre-defined skills.
@@
Affordance reasoning for robotic control. The psychologist James J. Gibson, along with Eleanor J. Gibson, introduced the concept of an affordance [11], which refers to the ability to perform a certain action with an object in a given environment. Much of the research related to affordances has concentrated on predicting how to interact with objects, as demonstrated by [47, 12, 3]. In the field of robot manipulation, keypoints are often used to provide compact information about the environment and the objects [6, 9, 48, 31, 33, 32, 38], representing the affordance in a structured way. Among these works, the most related one to ours is KETO [38], which predicts affordance and functional keypoints on the tool objects by learning from interactions with a trajectory optimization formulation as in Manuelli et al. [32]. In contrast to KETO [38], our keypoints selection procedure doesn’t require any model training. We design an automated process to annotate keypoints as visual marks on a 2D image, and leverage the broad knowledge from GPT-4V to select affordance and motion keypoints for manipulation.
@@ III Problem Statement
Our goal is to enable robots to perform manipulation tasks involving unseen objects and goals. Each task is described by a free-form language description $l$. To achieve success, the robot needs to interact with the environment across one or multiple stages. An example task is shown in Figure 1, where the robot is commanded by “Swipe the snack package off the table. Be careful with the eyeglasses on the table.”.
@@
We refer to the physical interaction at each stage as a subtask. These subtasks can include interactions with objects in hand (e.g., lifting up an object, opening a drawer, turning the faucet), interactions with environmental objects unattached to the robot (e.g., pushing an obstacle, pressing a button), and tool use which involves grasping a tool object to make contact with another object (e.g., poking with a stick, sweeping with a foam, cutting with a knife). In this example, the robot needs to perform two subtasks: First, put the eyeglasses, which are between the broom and the snack package, back into the case; and second, use the broom to sweep the trash. Since we do not assume the sequence of subtasks is given beforehand, the robot needs to decompose the task into a sequence of feasible subtasks based on the free-form language description $l$ and control the robot to solve each subtask.
@@
We consider a table-top manipulation setting that involves a robotic arm and RGBD cameras, as shown in Figure 2. At each time step $t$, an observation $s_{t}$ is received from the environment and an action $a_{t}$ is chosen to command the robot. The observation $s_{t}$ consists of the RGBD images captured by the camera sensors as well as the proprioception information. The action $a_{t}$ is defined in the Cartesian space.
@@
In this paper, we aim to solve such tasks by leveraging vision-language models (VLMs) with an effective visual prompting strategy. We use $\mathcal{M}$ to denote the VLM $\mathcal{M}$ which can take varying numbers of language and visual inputs in a specific order, and generate text responses accordingly. The responses of $\mathcal{M}$ can be controlled by designing different input prompts, which specify information such as the problem descriptions, the input and output formats, and in-context examples [17, 25, 28, 45]. Specifically, we designed text and visual prompts to enable $\mathcal{M}$ to generate desired text responses in a structured format, which can be easily parsed for downstream usage.
IV Marking Open-vocabulary Keypoint Affordances
We propose Marking Open-vocabulary Keypoint Affordances  (MOKA), an approach that leverages the emergent reasoning capability of Vision-Language Models (VLMs) to guide a low-level motion generator to solve unseen tasks specified by free-form language instructions. As shown in Fig. 2, MOKA uses a point-based affordance representation to connect the VLM’s prediction on 2D images with the robot’s motion in the physical world. To effectively employ the pre-trained VLM, MOKA converts the problem of affordance reasoning into a series of visual question-answering problems that are feasible for VLMs to solve.
@@ IV Marking Open-vocabulary Keypoint Affordances
In this section, we will start by introducing our point-based affordance representations and how they are used for generating motions. Then, we will describe a hierarchical framework that effectively prompts VLMs for affordance reasoning. After this, we will explain a novel visual prompting approach that lifts the VLM’s reasoning capabilities by annotating a set of marks (including candidate points, grids, and labels) on 2D images. Lastly, we will investigate two ways of bootstrapping the performance of our approach through interactions collected in the physical world.
IV-A Motion with Point-based Affordances

### v3

# MOKA: Open-World Robotic Manipulation through Mark-Based Visual Prompting

@@ Abstract
Open-world generalization requires robotic systems to have a profound understanding of the physical world and the user command to solve diverse and complex tasks. While the recent advancement in vision-language models (VLMs) has offered unprecedented opportunities to solve open-world problems, how to leverage their capabilities to control robots remains a grand challenge. In this paper, we introduce Marking Open-world Keypoint Affordances (MOKA), an approach that employs VLMs to solve robotic manipulation tasks specified by free-form language instructions. Central to our approach is a compact point-based representation of affordance, which bridges the VLM’s predictions on observed images and the robot’s actions in the physical world. By prompting the pre-trained VLM, our approach utilizes the VLM’s commonsense knowledge and concept understanding acquired from broad data sources to predict affordances and generate motions. To facilitate the VLM’s reasoning in zero-shot and few-shot manners, we propose a visual prompting technique that annotates marks on images, converting affordance reasoning into a series of visual question-answering problems that are solvable by the VLM. We further explore methods to enhance performance with robot experiences collected by MOKA through in-context learning and policy distillation. We evaluate and analyze MOKA’s performance on various table-top manipulation tasks including tool use, deformable body manipulation, and object rearrangement.
Figure 1: To solve manipulation tasks with unseen objects and goals, MOKA employs a VLM to generate motions through a point-based affordance representation (plotted as colorful dots on the images). By annotating marks (e.g., candidate points, grids, and captions) on the observed 2D image, MOKA converts motion generation into a series of visual question-answering problems that the VLM can solve.
@@ I Introduction
The pursuit of open-world generalization poses a significant challenge for robotic systems: Solving various tasks commanded by humans in complex and diverse environments requires a profound understanding of the physical world. An appealing prospect for handling this challenge is to employ large pre-trained models that encapsulate extensive prior knowledge from broad data and bringing it to bear on novel problems. Recent advances in large language models (LLMs) and vision-language models (VLMs) provide particularly promising tools in this regard, with their emergent and fast-growing conceptual understanding, commonsense knowledge, and reasoning abilities [10, 50, 51, 6, 9, 1, 52, 31, 53, 32]. However, existing large models pre-trained on Internet-scale data still lack the capabilities to understand 3D space, contact physics, and robotic control, not to mention the knowledge about the embodiment and environment dynamics in each specific scenario. This creates a large gap between the promising trend in computer vision and natural language processing and applying them to robotics. It remains an open question how such tools can guide a robotic system to solve manipulation tasks in the physical world.
@@
Recently, a growing body of research has been dedicated to utilizing pre-trained large-scale models for robotic control. To incorporate broad knowledge, these approaches either directly prompt or fine-tune the large models to generate plans [2, 22, 21, 7], rewards [38, 29, 40, 68], codes [34, 56, 61], etc. Despite the encouraging results they have demonstrated, these approaches are subject to notable limitations. Since the advances in LLMs precede VLMs, many previous approaches first process the raw sensory inputs to obtain the language description of the environment and then query LLMs to perform reasoning and planning in the language domain. However, relying solely on high-level language descriptions may overlook the nuanced visual details of environments and objects, which are vital for accurately completing tasks. Moreover, existing approaches usually require non-trivial effort in designing in-context examples [23, 34, 56] to ensure LLMs can produce desired predictions on similar tasks. As a result, the tasks that can be solved by these approaches are largely constrained by such manual efforts.
@@
In this work, we study how to effectively endow robots with the ability to solve novel manipulation tasks specified by free-form language instructions using VLMs. Our key insight is to find an intermediate affordance representation that connects the VLM’s prediction on images with the robot’s motion in the physical world. This affordance representation should satisfy two essential requirements. First, it should be feasible for the VLM to predict such representation based on the visual observation of the environment and the task description. Second, the representation should compactly capture the information about the relevant properties of the robot’s motion so that it can be easily executed on the robot.
@@
To this end, we propose Marking Open-world Keypoint Affordances (MOKA), an approach that employs VLMs for robotic manipulation through mark-based visual prompting. As shown in Fig. 1, MOKA leverages a compact affordance representation consisting of a set of keypoints and waypoints, defined on open sets of objects and tasks. This point-based affordance representation is then used to specify the desired motion for the robot to solve the task. To generate the motions given the free-form language descriptions, MOKA uses hierarchical visual prompting to convert the affordance reasoning problem into a series of visual question-answering problems. Drawing inspirations from recent advances in visual question-answering [63], we use mark-based visual prompting to enable the VLM to attend to the important visual cues in the observation image and further simplify the point generaton problem into multiple choice questions. As shown in the top-left part of Fig. 1, we plot candidate points on the image and query the VLM to select the optimal set of points that will result in the desired motion. The predicted keypoints and waypoints are used to specify a motion trajectory which can cover a wide range of manipulation skills such as picking, placing, pressing, tool use, etc.
@@
We demonstrate the effectiveness of MOKA on various manipulation tasks, with robustness across the variations of instructions, objects, and initial arrangements. We further use MOKA to collect successful trajectories in each task. The collected trajectories can be used as in-context examples to bootstrap the performance of VLM, and can also be leveraged as demonstrations to train a better student policy. Our experiments show that MOKA can achieve state-of-the-art performance on various evaluation tasks in a zero-shot manner, with consistent improvements using clean and intuitive in-context examples. To summarize, our contributions are:
@@
  • We propose a mark-based visual prompting approach that converts affordance reasoning into a series of visual question-answering problems.
  • We demonstrate that the proposed approach, MOKA, can effectively generate motions for open-world manipulation in zero-shot and few-shot manners.
@@ II Related Work
Figure 2: Overview of MOKA. We propose a hierarchical approach to prompt the VLM to perform affordance reasoning based on the free-form langauage instruction of the task and the visual observation of the environment. On the high level, we query the VLM to decompose the free-form language description of the task into a sequence of subtasks and summarize the subtask information. On the low level, the VLM is prompted to produce the keypoints and additional attributes for the affordance representation defined in Sec. IV-A.
@@
Large language models and vision-language models. Recent advances in several domains and applications have been greatly influenced by the substantial progress achieved through large language models (LLMs) and vision language models (VLMs) [10, 50, 51, 6, 9, 1, 52, 31, 53, 32]. While these models can already solve various tasks in a zero-shot manner [1], well-designed prompts still serve an important role in further eliciting more advanced capabilities. As demonstrated by Brown et al. [6], few-shot prompting can match or surpass the performance of fine-tuning on LLMs. Additionally, other prompting techniques [62, 69, 28, 67] have been proposed to improve LLMs’ reasoning capabilities. Although these methodologies may initially appear enigmatic, they have been empirically validated to consistently demonstrate scalability across various models [6, 9, 1, 66]. Apart from the language prompts, recent vision models [27, 70, 30] are capable of supporting visual prompts, including points, boxes, masks, and texts. Such visual prompts exhibit greater diversity in both form and content, harnessing vision-language models for various application scenarios like perception [65], image editing [55] and reasoning [63].
@@
Incorporating robust visual reasoning capabilities becomes imperative to build autonomous robots capable of making decisions in unstructured environments. Although the predictions by existing VLMs cannot seamlessly engage in zero-shot reasoning, they can still be effectively harnessed across diverse tasks. Prior work such as SoM [63] draws visual marks on the objects in an image using numbers and segmentation masks and demonstrates the mark-based visual prompting scheme unleashes the reasoning capabilities of GPT-4V, such as object counting and inferring spatial relationship. Inspired by SoM [63], we wleverage GPT-4V for robotic manipulation. We represent manipulation skills with a set of affordance points and motion waypoints. Instead of using object segmentation masks, we use keypoints and grid cells as visual prompts as shown in Fig. 1, and then query GPT-4V to choose the affordance points and motion waypoints from the visual marks, followed by executable point-based motion plans.
@@
Foundation models in robotics. Building upon the successes of foundation models, the field of robotics is currently witnessing a rising interest in utilizing LLMs and other foundation models in various applications. With large model capacities, transformers [60] and VLMs can be trained from scratch or fine-tuned to directly predict actions through imitation learning and reinforcement learning [12, 5, 24, 37, 45]. However, training large models for open-world scenarios would usually require extensive data and supervision. Pre-trained LLMs can also be used to generate high-level task plan in natural language [2, 22, 21, 7], executable programs [34, 56, 61] or value function [23, 35] for low-level behaviours, environmental reward and feedback for reinforcement learning [38, 29, 40, 68]. However, many of these approaches necessitate the conversion of both the task and the observations into the textual form. While this can be easily accomplished in simulated environments with ground-truth object states, tasks in the real world require the utilization of robust and precise perception modules. For performing open-ended navigation and manipulation in the real world [57, 7, 16], open-world VLMs [27, 70, 30, 36, 44, 33] are often used to extract visual scene information before converting the observation to textual form. However, the process of converting an image into text may result in the loss of important details, such as shape and geometric information. With the recent advancements in VLMs, it is now conceivable that the role of state estimation combined with language models can be replaced by a single, powerful VLM, such as GPT-4V, which is leveraged by MOKA.  Hu et al. [20] utilizes GPT-4V to get semantic language plans and convert each language plan to a set of pre-defined low-level skills. Different from Hu et al. [20], we represent affordances and motions using a set of points, and query GPT-4V to select the corresponding points (e.g. grasping point, function point) in each task, which can be directly translated to low-level point-based motions. While preserving the visual reasoning capabilities of VLMs, MOKA provides a framework with more general and flexible low-level motion, which doesn’t require prior knowledge about a specific task compared to pre-defined skills.
@@
Affordance reasoning for robotic control. Affordance [17] refers to the ability to perform a certain action with an object in a given environment. Much of the research on affordances focuses on understanding interactions with objects, as demonstrated by [58, 18, 3]. In the field of robot manipulation, keypoints and other representation forms are often used to provide compact information about the environment and the objects [8, 15, 59, 41, 43, 11, 42, 49, 4, 25, 26, 13, 14], representing the affordance in a structured way. Among these works, the most related one to ours is KETO [49], which predicts affordance and functional keypoints on the tool objects by learning from interactions with a trajectory optimization formulation as in Manuelli et al. [42]. In contrast to KETO [49], our keypoints selection procedure doesn’t require any model training. We design an automated process to annotate keypoints as visual marks on a 2D image and leverage the broad knowledge from GPT-4V to select affordance and motion keypoints for manipulation.
@@ III Problem Statement
Our goal is to enable robots to perform manipulation tasks involving unseen objects and goals. Each task is described by a free-form language instruction $l$. To achieve the success, the robot needs to interact with the environment across one or multiple stages. An example task is shown in Fig. 1, in which the robot is commanded to clean the table by the instruction “Wipe the trash off the table after moving away the eyeglasses.”.
@@
The physical interaction at each stage is referred to as a subtask. Each subtask can be an interaction with objects in hand (e.g., lifting up an object, opening a drawer, turning the faucet), an interaction with environmental objects unattached to the robot (e.g., pushing an obstacle, pressing a button), and tool use involving grasping a tool object to make contact with another object (e.g., poking with a stick, sweeping with a foam, cutting with a knife). In this example above, the robot needs to perform two subtasks: First, put the eyeglasses between the broom and the snack package back into the case; and second, use the broom to sweep the trash. Since we do not assume the sequence of subtasks to be given beforehand, the robot needs to decompose the task into a sequence of feasible subtasks based on the free-form language instruction $l$ and solve each subtask sequentially.
@@
We consider a table-top setting with a robotic arm and RGBD cameras, as shown in Fig. 2. At each time step $t$, an observation $s_{t}$ is received from the environment and an action $a_{t}$ is selected to command the robot. The observation $s_{t}$ consists of the RGBD images captured by the camera sensors as well as the proprioception of the robot. The action $a_{t}$ is defined as the 6-DoF gripper pose and the finger status.
@@
In this paper, we aim to solve such tasks by leveraging vision-language models (VLMs) with an effective visual prompting strategy. We use $\mathcal{M}$ to denote the VLM, which can take a list of language and visual inputs in a specific order, and generate text responses accordingly. The responses of $\mathcal{M}$ can be controlled by designing different input prompts, which specify information such as the problem descriptions, the input and output formats, and in-context examples [23, 34, 38, 56]. Specifically, we design text and visual prompts to enable $\mathcal{M}$ to generate desired text responses in a structured format, which can be parsed into point coordinates and other attributes to characterize the robot motion for the downstream usage.
IV Marking Open-world Keypoint Affordances
@@
We propose Marking Open-world Keypoint Affordances  (MOKA), an approach that leverages the emergent reasoning capability of Vision-Language Models (VLMs) to guide a motion planner to solve unseen tasks specified by free-form language instructions (see Fig. 2). In this section, we will start by introducing the point-based affordance representations and how they are used for generating motions. Then, we will describe a hierarchical framework that effectively prompts VLMs for affordance reasoning. After this, we will explain a novel visual prompting approach that lifts the VLM’s reasoning capabilities by annotating a set of marks, including candidate points, grids, and labels, on 2D images. Lastly, we will investigate two ways to bootstrap the performance of MOKAby utilizing successful experiences collected in the physical world.
IV-A Point-Based Affordance Representations